OpenStudy (anonymous):

If y = f ( x) is defined by the given equation, find the expression for f ( x). Given: 3x - 2y = 6 f ( x) = ?

OpenStudy (apoorvk):

rearrange the equation to write y solely in terms of x. just shifting and all!

OpenStudy (bahrom7893):

Do what apoorvk said: 3x-2y=6 -2y=6-3x y=(3/2)x-3

OpenStudy (bahrom7893):

f(x)=(3/2)x-3

OpenStudy (anonymous):

Thanks bahrom and apoorvk!